OpenStudy (anonymous):

x and y vary inversely, and x = 1/5 when y = 5. What is the value of k?

OpenStudy (anonymous):

x=k/y put x=1/5 and y=5 k=1

OpenStudy (anonymous):

Thank you! Could you help me on 2 two more please?

OpenStudy (anonymous):

use the method I used for inverse proportion x=k/y for direct proportion use x=ky